# LLaVA-PruMerge: Adaptive Token Reduction for Efficient Large Multimodal Models

### v5

@@ 5 Conclusion
Appendix A Appendix
A.1 Limitations and Future Work
Our exploration of LLaVA-PruMerge has two primary limitations. First, the visual token compression, while efficient, is not entirely lossless. This results in a marginal performance gap between the original model (LLaVA) and our optimized version LLaVA-PruMerge. We are committed to advancing our research towards achieving a fully lossless token compression algorithm, aiming to close these performance gaps. Second, the scope of our validation efforts is somewhat constrained by the computational resources typically available in academic settings. This limitation has precluded a comprehensive assessment of PruMerge’s applicability to larger-scale models, such as those envisioned in the LLaVA-Next [Liu et al., 2024] framework with a Yi-34B backbone. Future investigations will seek to extend our methodology to these more expansive model architectures, exploring its potential for generalization and broader impact.
A.2 Societal Impacts
In this study, we propose a technique that improves the efficiency of Large Multimodal Models (LMMs), making them more accessible. This approach helps to democratize LMMs by lowering deployment costs and hardware barriers, facilitating their use in edge computing. However, it does not mitigate the potential misuse of LMMs by malicious actors.

### v6

@@ 5 Conclusion
6 Acknowledgments
This research is supported by NSF IIS-2525840 (Yan Yan), CNS-2432534 (Yan Yan), ECCS-2514574 (Yan Yan), NIH 1RF1MH133764-01 (Yan Yan) Cisco Research unrestricted gift (Yan Yan), NSF IIS2404180 (Yong Jae Lee), Microsoft Accelerate Foundation Models Research Program (Yong Jae Lee), and Institute of Information & communications Technology Planning & Evaluation (IITP) grants funded by the Korea government (MSIT) (No. 2022-0-00871, Development of AI Autonomy and Knowledge Enhancement for AI Agent Collaboration, Yong Jae Lee) and (No. RS-2022-00187238, Development of Large Korean Language Model Technology for Efficient Pre-training, Yong Jae Lee). This article solely reflects opinions and conclusions of authors and not funding agencies.